OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P?

OpenStudy (anonymous):

Definition of a circle: (x-a)^2 + (y-b)^2 = R^2, with (a,b) the centre and R the Radius.

OpenStudy (anonymous):

(x-g)^2+(y-f)^2=8^2

OpenStudy (anonymous):

Where g=2 and f=-1

OpenStudy (anonymous):

Use the formula Substitute the values Get the equation Thats ur answer

OpenStudy (anonymous):

Cheers:)

OpenStudy (anonymous):

Find it out!!:)